write an equation for
with a slope of negative three-fourths and a y-intercept of -3
write an equation for
through (0,1) and with a slope of 1.5

Answer by nerdybill(7384)

write an equation for
with a slope of negative three-fourths and a y-intercept of -3
.
"slope-intercept" form is:
y = mx + b
plug in given values:
y = (-3/4)x - 3 (this is what they're looking for)
.
write an equation for
through (0,1) and with a slope of 1.5
plug above into "point-slope" form:
y - y1 = m(x - x1)
y - 1 = 1.5(x - 0)
y - 1 = 1.5x
y = 1.5x + 1